# Permutation twisted cohomology, remixed

### v2

@@ 4.2. The open cover
Given this observation, we generalize the open cover of $\operatorname{Spc}(\mathscr{K})$ presented in [BG25, Proposition 13.11] as follows. Let $\mathcal{B}(G)$ denote the subset of the canonical $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$ not induced from the trivial $\mathbb{R}G$-module $\mathbb{R}$, i.e. the effective endotrivials arising from irreducible real representation spheres, excluding $k[1]$. Define an open of $\operatorname{Spc}(\mathscr{K})$ by
@@
Note that in defining the opens $U(H)$, the canonical $\mathbb{Z}$-basis element $k[1]\in\mathcal{B}(G)$ would play no significant role, so we may exclude $k[1]$ from $\mathcal{B}(G)$ without issue.
@@
Let us recall the closed points of $\operatorname{Spc}(\mathscr{K})$. Recall that closed points are exactly the minimal primes for inclusion, and every prime contains a minimal prime. [BG25, Corollary 7.31] asserts that the minimal primes of $\operatorname{Spc}(\mathscr{K})$ are those explicitly of the form $\mathfrak{m}_{H}:=\mathscr{P}(H,0)$ for some subgroup $H$, where 0 denotes the unique closed point of $V_{{G}/\!\!/{H}}$. These closed points are precisely the kernel of the residue tt-functor $\mathbb{F}^{H}:=\operatorname{Res}_{1}^{{G}/\!\!/{H}}\circ\Psi^{H}=\Psi^{H}\circ\operatorname{Res}^{G}_{H}$ (see [BG25, Definition 7.26]).
The collection $\{U(H)\}_{H\leq G}$ is an open cover of $\operatorname{Spc}(\mathscr{K})$, with the closed point $\mathfrak{m}_{H}$ belonging to $U(H)$.
@@
Let $H$ be a subgroup of $G$ and $C$ be an effective endotrivial. Recall the residue tt-functor $\mathbb{F}^{H}\colon\operatorname{Res}^{{G}/\!\!/{H}}_{1}\circ\Psi^{H}\colon\mathscr{K}\to\operatorname{D}_{b}(k)$ of the closed point $\mathfrak{m}_{H}$. $\mathbb{F}^{H}(\iota^{H}_{C})$ is an isomorphism. Moreover, if $K$ is a subgroup of $G$ satisfying $h_{C}(K)\neq h_{C}(H)$, $\mathbb{F}^{K}(\iota^{H}_{C})$ is not an isomorphism.
In particular, the open $U(H)$ contains $\mathfrak{m}_{H}$, therefore the set of opens $\{U(H)\}_{H\leq G}$ is an open cover of $\operatorname{Spc}(\mathscr{K})$.
@@
The fact that $\mathbb{F}^{H}(\iota^{H}_{C})$ is an isomorphism follows immediately since $\Psi^{H}(\iota^{H}_{C})$ is an isomorphism. Moreover, if $h_{C}(K)\neq h_{C}(H)$, $\mathbb{F}^{K}(\iota^{H}_{C})$ cannot be an isomorphism since $k[h_{C}(H)]$ is not isomorphic to $\Psi^{K}(C)$ in $\operatorname{D}_{b}(k[{G}/\!\!/{K}])$.
It follows by conservativity that for any effective endotrivial $C$, $\iota^{H}_{C}$ is an isomorphism in $\mathscr{K}/\mathfrak{m}_{H}$ since $\mathbb{F}^{H}$ is the residue functor of $\mathfrak{m}_{H}$, so $U(H)$ contains $\mathfrak{m}_{H}$. Therefore by general tt-geometry, $\{U(H)\}_{H\leq G}$ is an open cover of $\operatorname{Spc}(\mathscr{K})$, as every prime specializes to some $\mathfrak{m}_{H}$, which are precisely the closed points of $\operatorname{Spc}(\mathscr{K}(G)$. ∎
We next show that if $K,H$ are non-conjugate subgroups of $G$, then $\mathfrak{m}_{K}\not\in U(H)$. This takes a bit more work. It follows from the previous proposition that if there exists a Borel-Smith function $f$ for which $f(K)\neq f(H)$, then $\mathfrak{m}_{K}\not\in U(H)$. However, this may not necessarily occur!
@@
This group has a GAP implementation of SmallGroup(32,43); it is the holomorph of $\mathbb{Z}/8\mathbb{Z}$. In this case, $C_{2}\times C_{2}$ acts on $D_{8}$ faithfully. $G$ has two nonconjugate subgroups isomorphic to $V_{4}$,
@@
Since $\operatorname{im}(\dim)=\operatorname{CF}_{b}(G)$, every Borel-Smith function $f$ satisfies $f(H)=f(K)$. The author thanks Math.SE user testaccount for this example.
@@
Let $H,K$ be non-conjugate subgroups of $G$ with the same order, let $L$ be the smallest subgroup of $G$ containing both a conjugate of $H$ and a conjugate of $K$, and let $C$ be an effective endotrivial for which $h_{C}(L)<h_{C}(H)$. Then $\mathbb{F}^{K}(\iota^{H}_{C})$ and $\mathbb{F}^{H}(\iota^{K}_{C})$ are 0.
@@
It suffices to assume $h_{C}(H)=h_{C}(K)$. From Proposition 4.1 and the construction of $\iota^{H}_{C}$ in Theorem 4.2, the image of $\iota^{H}_{C}$ in homological degree $h_{C}(H)$ is contained in an indecomposable permutation $kG$-module $M$ with stabilizer containing $H$ but strictly contained in $L$. Therefore, $\Psi^{K}(M)=0$, so the image of $\mathbb{F}^{K}(\iota^{H}_{C})$ in degree $h_{C}(H)$ is the zero map, as desired. An analogous argument demonstrates $\mathbb{F}^{H}(\iota^{K}_{C})$ is also the zero map. ∎
Corollary 4.11.
The prime $\mathfrak{m}_{H}$ is the only closed point of $\operatorname{Spc}(\mathscr{K})$ contained in the open $U(H)$.
@@
We previously showed in Proposition 4.7 that $\mathfrak{m}_{H}\in U(H)$. Since $U(H)$ is defined by iterating over the irreducible endotrivials, if $h_{C}(H)\neq h_{C}(K)$ for some irreducible endotrivial $C$, then $\mathfrak{m}_{K}\not\in U(H)$. Therefore, it suffices to consider the case where $G$ has a pair of indistinguishable subgroups $H,K$. If this occurs, clearly $H,K<G$, so $H,K<_{G}L\leq G$, where $L$ is the smallest subgroup of $G$ containing both a conjugate of $H$ and $K$. Since $|L|>|H|$, there exists a Borel-Smith function $f$ for which $f(L)\neq f(H)$ (see Remark 4.9). Therefore, there exists a canonical $\mathbb{Z}$-basis element $b$ of $\operatorname{CF}_{b}(G)$ which also satisfies $b(L)\neq b(H)$. Hence, there exists an element $C$ of the canonical $\mathbb{Z}$-basis $\mathcal{B}(G)\cup\{k[1]\}$ of $\operatorname{Pic}(\mathscr{K}(G))$ with $h_{C}(H)\neq h_{C}(L)$. The previous lemma implies $\mathbb{F}^{K}(\iota^{H}_{C})$ is not an isomorphism, and it follows by definition that $\mathfrak{m}_{K}\not\in U(H)$, as desired. ∎
Warning: Corollary 4.11 does not imply that the open $U(H)$ has a unique closed point viewed as a subspace of $\operatorname{Spc}(\mathscr{K})$. The open $U(H)$ has a minimal closed point if and only if $\mathscr{K}(G)|_{U(H)}\cong\mathscr{K}(G)/\mathfrak{m}_{H}$, i.e. $U(H)$ consists of all points specializing to $\mathfrak{m}_{H}$. In general this may not occur.
Remark 4.12.
Corollary 4.11 result does not hold for the open cover of [BG25, Section 13] for non elementary abelian $p$-groups. For instance, [BG25, Proposition 13.14] states the closed complement of the open $U^{\prime}(1)$ (we write $U^{\prime}(1)$ to denote the open $U_{G}(1)$ of [BG25, Proposition 13.14] in order to distinguish the covers apart) is the support of $\operatorname{kos}_{G}(F)$, where $F$ denotes the Frattini subgroup of $G$ (i.e. the intersection of all maximal subgroups of $G$). By [BG25, Corollary 7.17], we have $\mathfrak{m}_{H}\in\operatorname{supp}(\operatorname{kos}_{G}(F))$ if and only if $H\not\leq_{G}F$, hence $\mathfrak{m}_{H}\in U^{\prime}(1)$ if and only if $H\leq_{G}F$. Therefore, $U^{\prime}(1)$ contains a unique closed point of $\mathscr{K}(G)$ if and only if $F=1$ if and only if $G$ is an elementary abelian $p$-group.
Corollary 4.13.
@@
By construction of $U(H)$, every irreducible endotrivial corresponding to an element of the canonical $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$ is isomorphic to $k[h_{C}(H)]$ in $\mathscr{K}(G)|_{U(H)}$. As the set of irreducible endotrivials is a basis of $\operatorname{Pic}(\mathscr{K}(G))$, it follows that every endotrivial $C$ is isomorphic to $k[h_{C}(H)].$
@@
Corollary 4.13 does not imply the existence of a homomorphism $f:k[h_{C}(H)]\to C$ which is an isomorphism in $\mathscr{K}(G)|_{U(H)}$. Such homomorphisms may not exist in general for non-effective endotrivials. For example, let $p=2$ and $G=C_{2}$, then the two-term endotrivial $C:=k\xrightarrow{\operatorname{coaug}}kC_{2}$ with $k$ in homological degree 0 satisfies $\operatorname{Hom}_{\mathscr{K}(G)}(k,C)=0$, but we have a local isomorphism $k[0]\cong C$ in the localization $\mathscr{K}(G)|_{U(C_{2})}$.
We turn to our open $U(1).$ First, we need a lemma about faithful endotrivials, i.e. endotrivials that arise from a faithful real representation $V$. Equivalently, these are endotrivials whose h-marks are all 0 for any subgroup $H$ containing a nonzero normal subgroup of $G$, see [Mil24]. In this case, we recover a generalization of [BG25, Proposition 13.14].
@@
We have that $\ker(V)=1$, so by Proposition 4.1 if $n$ is the highest nonzero homological degree of $C$, $C_{n}\cong kG$. Now consider the complex $\operatorname{cone}(\iota^{1}_{C})^{*}$. After shifting, this complex satisfies $C_{1}\cong kG$ and $C_{0}\cong k$. Therefore, we are in the situation of [BG25, Corollary 3.20], and it follows that $\operatorname{cone}(\iota^{1}_{C})^{*}$, hence $\operatorname{cone}(\iota^{1}_{C})$, generates $\mathscr{K}_{ac}(G)$ as a tt-ideal, and $\operatorname{supp}(\mathscr{K}_{ac}(G))=\operatorname{supp}(\operatorname{cone}(\iota^{1}_{C}))$, as desired. ∎
@@
Analogously, if $V$ is a real representation with kernel $N\trianglelefteq G$, then it is a faithful real representation of $G/N$, and the associated endotrivial $C$ is a faithful endotrivial complex of $k[G/N]$-modules. It follows that in this case, $\operatorname{cone}(\iota^{1}_{C})=\operatorname{cone}(\iota^{N}_{C})$ (as they are the same map) and $\operatorname{supp}(\ker(\operatorname{Res}^{G}_{N}))=\operatorname{supp}(\operatorname{kos}_{G}(N))$, recovering [BG25, Lemma 13.2].
@@
Let $G$ be a finite $p$-group. The closed complement of the open $U(1)$ is the support of $\operatorname{kos}_{G}(1)$, i.e. the closed support of the tt-ideal $\mathscr{K}_{ac}(G)=\ker(\operatorname{Res}^{G}_{1})$. In particular, $U(1)$ is equal to the cohomological open $V_{G}=\operatorname{Spc}(\operatorname{D}_{b}(kG)).$
@@
Since $\mathcal{B}(G)\cup\{k[1]\}$ is a basis of $\operatorname{Pic}(\mathscr{K}(G))$, for any $\mathscr{P}\in U(1)$, all endotrivials are isomorphic to a shift of the tensor unit in $\mathscr{K}(G)/\mathscr{P}$. In particular, there exists a faithful real representation $V$ (for instance the regular representation $\mathbb{R}G$) and a corresponding endotrivial $C_{V}$, for which $\iota_{C_{V}}^{1}\colon k[h_{C_{V}}(1)]\to C_{V}$ is an isomorphism in $\mathscr{K}(G)/\mathscr{P}$ (in fact, an isomorphism in $\mathscr{K}|_{U(1)}$ by Corollary 4.13). Therefore, $\operatorname{open}(\operatorname{cone}(\iota^{1}_{C_{V}}))\supseteq U(1)$. By Proposition 4.15, we have $\operatorname{supp}(\operatorname{kos}_{G}(1))=\operatorname{supp}(\operatorname{cone}(\iota_{C_{V}}^{1}))\subseteq U(1)^{c}$.
@@
We have an tt-equivalence $\mathscr{K}(G)|_{U(1)}\cong\operatorname{D}_{b}(kG)$.
Proof.
This follows immediately from Proposition 4.17 and general tt-geometry. ∎
In particular, the open $U(1)$ has a unique closed point as a subspace of $\mathscr{K}(G)$. However, we stress that in general the other opens $U(H)$ need not have unique closed points.
5. The (re-)twisted cohomology ring
We are now ready to define the twisted cohomology ring for finite $p$-groups. When $G$ is a finite elementary abelian $p$-group, we recover the previously constructed twisted cohomology ring, since in this case the set of endotrivials $\{u_{N}\}_{N\in\mathcal{N}}$ as defined in [BG25, Example 12.1] and the shift of the tensor unit $k[1]$ form precisely the canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$.

### v3

@@ 4.2. An open cover of the spectrum
Given this observation, we generalize the open cover of $\operatorname{Spc}(\mathscr{K}(G))$ presented in [BG25, Proposition 13.11] as follows. Let $\mathcal{B}(G)$ denote the subset of the canonical $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$ not induced from the trivial $\mathbb{R}G$-module $\mathbb{R}$, i.e., the effective endotrivials arising from irreducible real representation spheres, excluding $k[1]$. For any subgroup $H\leq G$, we define an open subset of $\operatorname{Spc}(\mathscr{K}(G))$ via
@@
Since $\iota^{H}_{C}=\iota^{{}^{g}H}_{C}$ for any $g\in G$, $U(H)=U({}^{g}H)$. Moreover, $\operatorname{open}(\iota_{k[1]}^{H})=\operatorname{Spc}(\mathscr{K}(G))$ for any subgroup $H\leq G$. Therefore, there is no harm in excluding $k[1]=C_{\mathbb{R}}$ from $\mathcal{B}(G)$; this exclusion is mostly for notation and indexing purposes.
@@
Let us recall the closed points of $\operatorname{Spc}(\mathscr{K}(G))$, i.e., the minimal primes with respect to inclusion. [BG25, Corollary 7.31] asserts that the minimal primes of $\operatorname{Spc}(\mathscr{K}(G))$ are those $\mathscr{M}(H):=\mathscr{P}(H,0)$ for some subgroup $H$, where 0 denotes the unique closed point of $V_{{G}/\!\!/{H}}$. The closed point $\mathscr{M}(H)$ is precisely the kernel of the residue tt-functor $\mathbb{F}^{H}:=\operatorname{Res}_{1}^{{G}/\!\!/{H}}\circ\Psi^{H}=\Psi^{H}\circ\operatorname{Res}^{G}_{H}$, see [BG25, Definition 7.26], and $\mathcal{M}(H)$ is also the kernel of $\hat{\Psi}^{H}$.
The collection $\{U(H)\}_{H\leq G}$ is an open cover of $\operatorname{Spc}(\mathscr{K}(G))$, with the closed point $\mathscr{M}(H)$ belonging to $U(H)$.
@@
Let $H$ be a subgroup of $G$ and $C$ be an effective endotrivial. Recall the residue tt-functor $\mathbb{F}^{H}\colon\operatorname{Res}^{{G}/\!\!/{H}}_{1}\circ\Psi^{H}\colon\mathscr{K}(G)\to\operatorname{D}_{b}(k)$ of the closed point $\mathscr{M}(H)$. $\mathbb{F}^{H}(\iota^{H}_{C})$ is an isomorphism. Moreover, if $K$ is a subgroup of $G$ satisfying $h_{C}(K)\neq h_{C}(H)$, $\mathbb{F}^{K}(\iota^{H}_{C})$ is not an isomorphism.
In particular, the open $U(H)$ contains $\mathscr{M}(H)$, therefore the set of opens $\{U(H)\}_{H\leq G}$ is an open cover of $\operatorname{Spc}(\mathscr{K}(G))$.
@@
The fact that $\mathbb{F}^{H}(\iota^{H}_{C})$ is an isomorphism follows immediately since $\Psi^{H}(\iota^{H}_{C})$ is a quasi-isomorphism by construction. Moreover, if $h_{C}(K)\neq h_{C}(H)$, $\mathbb{F}^{K}(\iota^{H}_{C})$ cannot be an isomorphism since $k[h_{C}(H)]$ is not isomorphic to $\Psi^{K}(C)$ in $\operatorname{D}_{b}(k[{G}/\!\!/{K}])$.
It follows by conservativity that for any effective endotrivial $C$, $\iota^{H}_{C}$ is an isomorphism in $\mathscr{K}/\mathscr{M}(H)$, as $\mathbb{F}^{H}$ is the residue functor of $\mathscr{M}(H)$, so $U(H)$ contains $\mathscr{M}(H)$. Therefore by general tt-geometry, $\{U(H)\}_{H\leq G}$ is an open cover of $\operatorname{Spc}(\mathscr{K}(G))$, as every prime specializes to some $\mathscr{M}(H)$, which are precisely the closed points of $\operatorname{Spc}(\mathscr{K}(G)$. ∎
We next show that if $K,H$ are non-conjugate subgroups of $G$, then $\mathscr{M}(K)\not\in U(H)$. This takes a bit more work. It follows from the previous proposition that if there exists a Borel-Smith function $f$ for which $f(K)\neq f(H)$, then $\mathscr{M}(K)\not\in U(H)$. However, this may not necessarily occur! We take a brief detour to explore this.
@@
This group has a GAP implementation of SmallGroup(32,43); it is the holomorph of $\mathbb{Z}/8\mathbb{Z}$. In this case, $C_{2}\times C_{2}$ acts on $D_{8}$ faithfully. $G$ has two nonconjugate subgroups isomorphic to $V_{4}=C_{2}\times C_{2}$,
@@
Since $\operatorname{im}(\dim)=\operatorname{CF}_{b}(G)$, every Borel-Smith function $f$ satisfies $f(H)=f(K)$. We thank Math.SE user testaccount for finding and sharing this example [TES24].
@@
Let $H,K$ be non-conjugate subgroups of $G$ with the same order, let $L$ be the smallest subgroup of $G$ for which $H,K\leq_{G}L$, and let $C$ be an effective endotrivial for which $h_{C}(L)<h_{C}(H)$. Then $\mathbb{F}^{K}(\iota^{H}_{C})=0$ and $\mathbb{F}^{H}(\iota^{K}_{C})=0$.
@@
It suffices to assume $h_{C}(H)=h_{C}(K)$. From 4.1 and the construction of $\iota^{H}_{C}$ in Theorem 4.3, there exists a direct sum decomposition of $C_{h_{C}(H)}$ such that the image of $\iota^{H}_{C}$ in homological degree $h_{C}(H)$ is contained in an indecomposable permutation $kG$-module summand $M$, with stabilizer containing $H$ but strictly contained in $L$. Therefore, by minimality of $L$, $\Psi^{K}(M)=0$, so $\mathbb{F}^{K}(\iota^{H}_{C})=0$, as desired. An analogous argument demonstrates $\mathbb{F}^{H}(\iota^{K}_{C})=0$. ∎
Putting everything together, we show that our open cover stratifies the closed points.
Theorem 4.12.
The prime $\mathscr{M}(H)$ is the only closed point of $\operatorname{Spc}(\mathscr{K}(G))$ contained in the open $U(H)$.
@@
We previously showed in 4.8 that $\mathscr{M}(H)\in U(H)$. Since $U(H)$ is defined by iterating over the irreducible endotrivials, if $h_{C}(H)\neq h_{C}(K)$ for some irreducible endotrivial $C$, then $\mathscr{M}(K)\not\in U(H)$. Therefore, it suffices to consider the case where $G$ has a pair of indistinguishable subgroups $H,K$. If this occurs, clearly $H,K<G$, so $H,K<_{G}L\leq G$, where $L$ is the smallest subgroup of $G$ containing both a conjugate of $H$ and $K$. Since $|L|>|H|$, there exists a Borel-Smith function $f$ for which $f(L)\neq f(H)$ (see Remark 4.10). Therefore, there exists a canonical $\mathbb{Z}$-basis element $b$ of $\operatorname{CF}_{b}(G)$ which also satisfies $b(L)\neq b(H)$. Hence, there exists an element $C$ of the canonical $\mathbb{Z}$-basis $\mathcal{B}(G)\cup\{k[1]\}$ of $\operatorname{Pic}(\mathscr{K}(G))$ with $h_{C}(H)\neq h_{C}(L)$. The previous lemma implies $\mathbb{F}^{K}(\iota^{H}_{C})$ is not an isomorphism, so by definition, $\mathscr{M}(K)\not\in U(H)$, as desired. ∎
Warning: Theorem 4.12 does not imply that the open $U(H)$ has a unique closed point, as a subspace of $\operatorname{Spc}(\mathscr{K}(G))$. The open $U(H)$ has a minimal closed point if and only if $U(H)$ consists of all points specializing to $\mathscr{M}(H)$. In general this may not occur; this can already be observed for elementary abelian $p$-groups.
As a result, we conclude that all endotrivials for $p$-groups are line bundles.
Theorem 4.13.
@@
By construction of $U(H)$ and the forerunners, every irreducible endotrivial corresponding to an element of the canonical $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$ is isomorphic to $k[h_{C}(H)]$ in $\mathscr{K}(G)|_{U(H)}$. As the set of irreducible endotrivials is a $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$, it follows that every endotrivial $C\in\operatorname{Pic}(\mathscr{K}(G))$ satisfies $C\cong k[h_{C}(H)]$ in $\mathscr{K}(G)|_{U(H)}$. ∎
@@
Given an arbitrary endotrivial $C$, Theorem 4.13 does not imply the existence of an $H$-forerunner, i.e., a homomorphism $f:k[h_{C}(H)]\to C$ which descends to an isomorphism in $\mathscr{K}(G)|_{U(H)}$. Such homomorphisms may not exist in general for non-effective endotrivials. For example, let $p=2$ and $G=C_{2}$, then the two-term endotrivial $C:=k\xrightarrow{\operatorname{coaug}}kC_{2}$ with $k$ in homological degree 0 satisfies $\operatorname{Hom}_{\mathscr{K}(G)}(k,C)=0$, but we have a local isomorphism $k[0]\cong C$ in the localization $\mathscr{K}(G)|_{U(C_{2})}$.
Definition 4.15.
We recall Balmer–Gallauer’s Koszul complexes: given a subgroup $H\leq G$, set
$\operatorname{kos}_{H}(G):=\operatorname{Ten}^{G}_{H}(0\to k\xrightarrow{\operatorname{id}}k\to 0),$
where $\operatorname{Ten}^{G}_{H}$ denotes tensor induction of chain complexes, see [BEN98b, Section 4.1]. Note that while tensor induction is monoidal, it does not preserve contractibility, and hence is not functorial for $\mathscr{K}(G)$. One has that the tt-ideal $\ker(\operatorname{Res}^{G}_{H})$ is generated by $\operatorname{kos}_{G}(H)$ by [BG25, Proposition 3.21]. In particular, $\operatorname{supp}(\operatorname{kos}_{G}(H))=\operatorname{supp}(\ker(\operatorname{Res}^{G}_{H}))$.
Remark 4.16.
Theorem 4.12 result does not hold for the open cover of [BG25, Section 13] for non elementary abelian $p$-groups, demonstrating that to stratify $\operatorname{Spc}(\mathscr{K}(G))$, one needs to consider more endotrivials than just the collection of $u_{N}$s. For instance, [BG25, Proposition 13.14] states the closed complement of the open $U^{\prime}(1)$ (writing $U^{\prime}(1)$ to denote the open $U_{G}(1)$ of [BG25, Proposition 13.14], in order to distinguish the open covers apart) is the support of $\operatorname{kos}_{G}(F)$, where $F$ denotes the Frattini subgroup of $G$ (i.e., the intersection of all maximal subgroups of $G$). By [BG25, Corollary 7.17], we have $\mathscr{M}(H)\in\operatorname{supp}(\operatorname{kos}_{G}(F))$ if and only if $H\not\leq_{G}F$, hence $\mathscr{M}(H)\in U^{\prime}(1)$ if and only if $H\leq_{G}F$. Therefore, $U^{\prime}(1)$ contains a unique closed point of $\mathscr{K}(G)$ if and only if $F=1$ if and only if $G$ is an elementary abelian $p$-group.
We turn to our open $U(1).$ Note that for any effective endotrivial $C$, $\iota^{1}_{C}$ is unique up to scalar since $C_{h_{C}(1)}$ is indecomposable by 4.1.
First, we need a lemma about faithful endotrivials, i.e., endotrivials that arise from a faithful real representation $V$. Equivalently, these are endotrivials whose h-marks are all 0 for any subgroup $H$ containing a nonzero normal subgroup of $G$, see [MIL24]. In this case, we recover a generalization of [BG25, Proposition 13.14].
@@
We have that $\ker(V)=1$, so by 4.1 if $n$ is the highest nonzero homological degree of $C$, $C_{n}\cong kG$. Now consider the dual complex $\operatorname{cone}(\iota^{1}_{C})^{*}$. After shifting, this complex satisfies $C_{1}\cong kG$ and $C_{0}\cong k$. Therefore, we are in the situation of [BG25, Corollary 3.20], so $\operatorname{cone}(\iota^{1}_{C})^{*}$, hence $\operatorname{cone}(\iota^{1}_{C})$, generates $\mathscr{K}_{ac}(G)$ as a tt-ideal, and $\operatorname{supp}(\mathscr{K}_{ac}(G))=\operatorname{supp}(\operatorname{cone}(\iota^{1}_{C}))$, as desired. ∎
@@
Analogously, if $V$ is a real representation with kernel $N\trianglelefteq G$, then it is a faithful real representation of $G/N$, and the associated endotrivial $C_{V}$ is a faithful endotrivial complex of $k[G/N]$-modules. In this case, $\operatorname{cone}(\iota^{1}_{C_{V}})=\operatorname{cone}(\iota^{N}_{C_{V}})$ (as they are the same map) and
$\operatorname{supp}(\ker(\operatorname{Res}^{G}_{N}))=\operatorname{supp}(\operatorname{kos}_{G}(N))=\operatorname{supp}(\operatorname{Inf}^{G}_{G/N}(\operatorname{kos}_{G/N}(N/N))),$
with the rightmost equality implied by [BG25, Lemma 3.17]. Therefore we have
$\operatorname{supp}(\iota_{C_{V}}^{N})=\operatorname{supp}(\ker(\operatorname{Res}^{G}_{N})),$
generalizing [BG25, Lemma 13.2(b)] to arbitrary endotrivials.
@@
Let $G$ be a finite $p$-group. The closed complement of the open $U(1)$ is the support of $\operatorname{kos}_{G}(1)$, i.e., the closed support of the tt-ideal $\mathscr{K}_{ac}(G)=\ker(\operatorname{Res}^{G}_{1})$. In particular, $U(1)$ is equal to the cohomological open $V_{G}=\operatorname{Spc}(\operatorname{D}_{b}(kG)).$
@@
Since $\mathcal{B}(G)\cup\{k[1]\}$ is a basis of $\operatorname{Pic}(\mathscr{K}(G))$, for any $\mathscr{P}\in U(1)$, all endotrivials are isomorphic to a shift of the tensor unit in $\mathscr{K}(G)/\mathscr{P}$. In particular, there exists a faithful real representation $V$ (for instance the regular representation $\mathbb{R}G$) and a corresponding endotrivial $C_{V}$, for which $\iota_{C_{V}}^{1}\colon k[h_{C_{V}}(1)]\to C_{V}$ is an isomorphism in $\mathscr{K}(G)/\mathscr{P}$ (in fact, an isomorphism in $\mathscr{K}(G)|_{U(1)}$ by Theorem 4.13). Therefore, $\operatorname{open}(\operatorname{cone}(\iota^{1}_{C_{V}}))\supseteq U(1)$. By 4.17, we have $\operatorname{supp}(\operatorname{kos}_{G}(1))=\operatorname{supp}(\operatorname{cone}(\iota_{C_{V}}^{1}))\subseteq U(1)^{c}$.
@@
The following corollary is immediate by general tt-geometry, since $\mathscr{K}(G)$ is rigid.
@@
We have a tt-equivalence $\mathscr{K}(G)|_{U(1)}\cong\operatorname{D}_{b}(kG)$.
In particular, the open $U(1)$ has a unique closed point as a subspace of $\mathscr{K}(G)$. However, we stress that the other opens $U(H)$ need not have unique closed points. In particular, the spaces $\operatorname{Spc}(\mathscr{K}(G)/\mathscr{M}(H))\subseteq\operatorname{Spc}(\mathscr{K}(G)|_{U(H)})$ need not be homeomorphic, as the former has a unique closed point, but the latter may not.
5. The remixed twisted cohomology ring
We introduce the twisted cohomology ring for finite $p$-groups.